OpenStudy (anonymous):

Multiply and simplify (3w-1)(4w-3)

OpenStudy (anonymous):

\[(3w - 1)(4w - 3) \implies 3w(4w - 3) - 1(4w - 3) \implies 12w^2 - 9w - 4w + 3\]\[\implies ?\]

OpenStudy (anonymous):

12w^2 - 5w + 3

OpenStudy (anonymous):

\[\large (a -b)(c - d) = ac - ad - bc +bd\]

OpenStudy (anonymous):

-9 - 4 = ??

OpenStudy (anonymous):

Not quite. -9w - 4w is basically this: -(9w + 4w) = ?

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

So as entire what will you get now??

OpenStudy (anonymous):

Yup. So it is: \[(3w - 1)(4w - 3) = 12w^2 - 13w + 3\]